Find the integral surface of the partial differential equation (x-y)y^2p+(y-x)x2q=(x^2+y^2) which passes through xz= a^3,y=0


1
Expert's answer
2019-09-10T11:00:21-0400
"(x-y)y^2p+(y-x)x^2q=x^2+y^2"

Auxiliary equations


"{dx \over (x-y)y^2}={dy \over (y-x)x^2}={dz\over (x^2+y^2)}"

"{dx \over (x-y)y^2}={dy \over (y-x)x^2}""x^2dx=-y^2dy""x^3+y^3=C_1"

"{dx-dy \over xy^2-y^3-yx^2+x^3}={dz \over x^2+y^2}""{d(x-y) \over (x-y)(x^2+y^2)}={dz \over x^2+y^2}"

"{d(x-y) \over x-y}={dz \over 1}""\ln |x-y|=z+\ln C_2""|x-y|=C_2e^z""F(x^3+y^2, |x-y|e^{-z})=0"


For "xz=a^3, y=0"


"x^3=C_1""|x|=C_2e^{a^3\/x}"

The integral surface 


"x=C"